# Blending 3D Geometry and Machine Learning for Multi-View Stereopsis

### v1

@@ Abstract
Traditional multi-view stereo (MVS) methods primarily depend on photometric and geometric consistency constraints. In contrast, modern learning-based algorithms often rely on the plane sweep algorithm to infer 3D geometry, applying explicit geometric consistency (GC) checks only as a post-processing step, with no impact on the learning process itself. In this work, we introduce GC-MVSNet++, a novel approach that actively enforces geometric consistency of reference view depth maps across multiple source views (multi-view) and at various scales (multi-scale) during the learning phase (see 1). This integrated GC check significantly accelerates the learning process by directly penalizing geometrically inconsistent pixels, effectively halving the number of training iterations compared to other MVS methods. Furthermore, we introduce a densely connected cost regularization network with two distinct block designs—simple and feature-dense—optimized to harness dense feature connections for enhanced regularization. Extensive experiments demonstrate that our approach achieves a new state-of-the-art on the DTU and BlendedMVS datasets secure second place on the Tanks and Temples benchmark. To our knowledge, GC-MVSNet++ is the first method to enforce multi-view, multi-scale supervised geometric consistency during learning. Our code is available at https://github.com/vkvats/GC-MVSNet-PlusPlus
Keywords 3D/stereo scene analysis  $\cdot$ Vision and Scene Understanding  $\cdot$ Stereo/Multi-View Stereo  $\cdot$ Machine Learning
Acknowledgement: This work was supported by the Electronics and Telecommunications Research Institute (ETRI) grant funded by the Korean government. [24ZC1100, The research of the fundamental media·contents technologies].
@@ 1 Introduction
Traditional multi-view stereo (MVS) methods, such as Gipuma [1], Furu [2], COLMAP [3], and Tola [4], primarily focused on addressing photometric and geometric consistency constraints across multiple views. In contrast, recent machine learning-based MVS approaches [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16] have used deep networks to extract feature maps and combine them into 3D cost volumes, capturing subtle similarities between features. These advances—ranging from multi-level feature extraction [5, 10] and attention-driven feature matching [5, 17] to refined cost-volume creation  [5, 6, 8, 17] and innovative loss formulations [5, 6, 13]—have significantly improved the fidelity of depth estimates and point cloud reconstructions..
@@
This formulation of the loss function offers richer geometric cues that significantly expedite model learning. Our extensive experiments reveal that GC-MVSNet++ nearly halves the number of training iterations compared to other contemporary models [5, 6, 10, 13, 15]. Our approach not only sets a new benchmark for accuracy on the DTU [18] and BlendedMVS [19] datasets, but also secures the second position on the Tanks and Temples [20] benchmark. GC-MVSNet++ is novel in its use of multi-view, multi-scale geometric consistency checks during training. Extensive ablation experiments further underscore the efficacy of our proposed method. To summarize:
@@
    The versatility of our module allows it to be seamlessly integrated into various MVS pipelines to boost geometric consistency during training.
@@
Figure 1: Our multi-view, multi-scale geometric consistency process. It explicitly models the geometric consistency of the estimated depth map across multiple source views during training at all three stages – coarse, intermediate, and refine. At each stage, the ref. depth estimate map undergo forward-backward-reprojection (see Alg. 2) across $N$ src. views to generate the GC-penalty map, which is then multiplied with per-pixel depth error. This approach enables the model to learn geometric constraints more quickly and accurately, leading to improved reconstruction quality during inference.
@@ 2 Related Work
Furukawa and Ponce [2] propose a taxonomy that classifies MVS methods into four primary scene representations: volumetric fields [21, 22, 23, 24], point clouds [9, 25], 3D meshes [26], and depth maps [1, 3, 5, 6, 10, 12, 14, 15, 16, 27]. Depth map-based methods can be further divided into traditional techniques that rely on feature detection and geometric constraint solving [1, 2, 3, 27], and learning-based approaches [5, 6, 10, 12, 14, 15, 16], which have gained significant popularity in recent years.
@@
While these methods enhance multi-stage MVS pipelines by refining specific components, they generally lack explicit integration of multi-view geometric cues during the learning process. Consequently, these models depend on the limited geometric information from multiple source views and the cost function formulation during training. Xu and Tao [14] address this limitation with a multi-scale geometric consistency-guided MVS approach that uses multi-hypothesis joint view selection to leverage structured region information for improved candidate hypothesis sampling. They argue that upsampled depth maps from source images can impose geometric constraints on estimates and utilize reprojection error [3, 33] to assess consistency. In contrast, our method employs forward-backward reprojection across multiple source views to directly evaluate and enforce geometric consistency in depth estimates, generating per-pixel penalties for geometrically inconsistent pixels.
@@
3 Methodology
MVS methods take $N$ views as input, including a reference image $I_{0}\in\mathbb{R}^{H\times W\times 3}$ and its paired ($N-1$) source view images $I^{N-1}_{i=1}$, along with the corresponding camera parameters $c_{0},...,c_{N}$, and then to estimate the reference view depth map ($D_{0}$) as the output.
3.1 Network Overview
The architecture of our approach Geometric Consistency MVSNet++, abbreviated as GC-MVSNet++, is shown in 2. We use a deformable convolution-based [41] feature pyramid network (FPN) [42] architecture (Sec. 3.4) to extract features from input images in a coarse-to-fine manner in three stages. At only the coarse stage, we apply a Global Feature Aggregation (GFA) module with linear attention [32, 5] to leverage global context information within and between reference and source view features. At each stage, we build a correlation-based cost volume of shape $H^{\prime}\times W^{\prime}\times D^{\prime}_{h}\times 1$ using feature maps of shape $N\times H^{\prime}\times W^{\prime}\times C$, where $H^{\prime}$, $W^{\prime}$, and $C$ denote the height, width, and number of channels of a given stage, and $D^{\prime}_{h}$ is the number of depth hypotheses at the corresponding stage. The cost volume is regularized with the proposed dense-connected CostRegNet. We use a winner-takes-all strategy to estimate the depth map $D_{0}$ at each stage.
We employ the GC module at each stage. It checks the geometric consistency of each pixel in $D_{0}$ across $M$ source views and generates $\xi_{p}$ (Sec. 3.2), a pixel-wise factor that is multiplied with the per-pixel depth error ($\xi_{d}$), calculated using a cross-entropy function. It penalizes each pixel in $D_{0}$ for its inconsistency across $M$ source views to accelerate geometric cues learning during training. TransMVSNet [5] trained with cross-entropy loss (GC-MVS-base) is our baseline; see Table 6 for different stages of GC-MVSNet++.
3.2 Multi-View Geometric Consistency Module
GC-MVSNet++ estimates reference depth maps at three stages with different resolutions. At each stage, the GC module takes $D_{0}$, $M$ source view ground truths $D^{gt}_{1},...D^{gt}_{M}$, and their camera parameters $c_{0},...c_{M}$ as input (see Alg. 1). The GC module is then initialized with a geometric inconsistency mask sum (or $mask\_sum$) of zero at each stage. This mask sum accumulates the inconsistency of each pixel across the $M$ source views. For each source view, the GC module performs forward-backward reprojection of $D_{0}$ to generate the penalty and then adds it to the mask sum.
Algorithm 1 Geometric Consistency Check Algorithm
@@ 3.2 Multi-View Geometric Consistency Module
  Output: $per\_pixel\_penalty$
  Require: $M\geq N$
@@
  $D\leftarrow\{D_{1}^{gt},\ldots,D_{M}^{gt}\}$
  $c\leftarrow\{c_{1}^{gt},\ldots,c_{M}^{gt}\}$
@@
     $D^{\prime\prime}_{P^{\prime\prime}_{0}},P^{\prime\prime}_{0}\leftarrow FBR(D_{0},c_{0},D_{i}^{gt},c_{i}^{gt})$
     $PDE\leftarrow\lVert P_{0}-P^{\prime\prime}_{0}\rVert_{2}$
     $RDD\leftarrow\frac{1}{D_{0}}\lVert D^{\prime\prime}_{P^{\prime\prime}_{0}}-D_{0}\rVert_{1}$
@@
  end for
  $per\_pixel\_penalty\leftarrow 1+mask\_sum/M$
Algorithm 2 Forward Backward Reprojection (FBR)
@@
  $D_{(R\rightarrow S)}\leftarrow K_{S}\cdot E_{S}\cdot E_{R}^{-1}\cdot K_{R}^{-1}\cdot D_{0}$
@@
  $D_{S_{remap}}\leftarrow REMAP(D_{i}^{gt},X_{D_{(R\rightarrow S)}},Y_{D_{(R\rightarrow S)}})$
  $D^{\prime\prime}_{P^{\prime\prime}_{0}}\leftarrow K_{R}\cdot E_{R}\cdot E_{S}^{-1}\cdot K_{S}^{-1}\cdot D_{S_{remap}}$
@@
For each stage, we generate two binary masks of inconsistent pixels, $PDE_{mask}$ and $RDD_{mask}$, by applying thresholds $D_{pixel}$ and $D_{depth}$, and then take a logical-OR of the two to produce a single mask of inconsistent pixels. These inconsistent pixels are assigned a value $1$ and all other pixels, including the consistent and the out-of-scope pixels, are assigned 00 to form a penalty mask. This penalty mask is then added to the mask sum (Alg. 1), which accumulates the penalty mask for each of the $M$ source views to generate a final mask sum with values $\in[0,M]$. Each pixel value indicates the number of inconsistencies of the pixel across the $M$ source views.
From this mask sum, we then generate the inconsistency penalty $\xi_{p}$ for each pixel. Our initial approach generated $\xi_{p}$ by dividing the mask sum by $M$ to normalize within the $[0,1]$. However, we found that using $\xi_{p}$ itself for element-wise multiplication reduces the contribution of perfectly consistent (zero inconsistency) pixels to zero, preventing further improvement of such pixels. To avoid this, we add $1$ so that elements of $\xi_{p}$ are in $[1,2]$. A reference view binary mask is applied on initial $\xi_{p}$ to generate the final $\xi_{p}$, as shown in Fig. 3.
Occlusion and its impact
In multi-view stereo, occluded pixels naturally occur as 3D points often remain invisible from certain views. These hidden pixels significantly affect geometric constraints, leading to the penalization of reference view pixels corresponding to occluded 3D points as inconsistent. To prevent these occluded pixels from disproportionately impacting geometric consistency losses, careful management is essential.
@@ 3.3.1 Dense Connectivity
Let us assume $H_{k}(.)$ denotes the feed-forward 2D convolution operation in a traditional 2D convolutional neural network (CNN), which feeds the output feature map $x_{k-1}$ of the $(k-1)^{th}$ layer to the input of the next layer, $x_{k}=H_{k-1}(x_{k-1})$. To bypass the non-linear transformation, ResNet [47] additionally adds skip-connections, $x_{k}=H_{k-1}(x_{k-1})+x_{k-1}$. DenseNet [46] further improves the information flow between layers by adding a denser connectivity pattern with direct connections from any layer to all its subsequent layers. More precisely, the $k^{th}$ layer receives the concatenation of feature maps [$x_{0},x_{1},...x_{k-1}$] from all the layers preceding it,
@@ 3.3.3 Decoder Dense Blocks
$\begin{split}x_{out}&=concat\left[x^{\prime}_{0},x^{\prime}_{1},...,x^{\prime}_{i-1}\right],\end{split}$ (2)
@@ 3.4 Deformable Feature Extraction Network
Unlike most MVS methods [5, 6, 10, 13, 51, 52] that use batch normalization [53] during training, we use group normalization throughout our network. As observed in [53], batch normalization provides more consistent and stable training with large batch sizes, but it is inconsistent and has a degrading effect on training with smaller batches. MVS methods are restricted to very small batch sizes, often between $1-4$, due to large memory requirements by the 3D-regularization network. Thus, we replaced batch normalization with group normalization layers [48] of group size $4$ across the network. Group normalization performs normalization across a number of channels that is independent of the number of examples in a batch [48]. We also implement weight standardization [49] for all layers (including 3D-convolutional layers) in the network. With these modifications, we achieve stable and reproducible training.
@@ 3.6 Cost Function
Table 1: Quantitative results on DTU and BlendedMVS. Accuracy (Acc), completeness (comp) and overall are in $mm$. We follow Darmon et al. [57] for BlendedMVS evaluation. Bold and underline represents first and second place, respectively.

### v2

@@ Abstract
Traditional multi-view stereo (MVS) methods primarily depend on photometric and geometric consistency constraints. In contrast, modern learning-based algorithms often rely on the plane sweep algorithm to infer 3D geometry, applying explicit geometric consistency (GC) checks only as a post-processing step, with no impact on the learning process itself. In this work, we introduce GC-MVSNet++, a novel approach that actively enforces geometric consistency of reference view depth maps across multiple source views (multi-view) and at various scales (multi-scale) during the learning phase (see Fig. 1). This integrated GC check significantly accelerates the learning process by directly penalizing geometrically inconsistent pixels, effectively halving the number of training iterations compared to other MVS methods. Furthermore, we introduce a densely connected cost regularization network with two distinct block designs—simple and feature-dense—optimized to harness dense feature connections for enhanced regularization. Extensive experiments demonstrate that our approach achieves a new state-of-the-art on the BlendedMVS dataset, and competitive performance on the DTU and Tanks and Temples benchmark. To our knowledge, GC-MVSNet++ is among the few approaches that enforce supervised geometric consistency across multiple views and at multiple scales during training. Our code is available at https://github.com/vkvats/GC-MVSNet-PlusPlus
keywords:
3D/stereo scene analysis , Vision and Scene Understanding , Stereo , Multi-View Stereo , Machine Learning , 3D Geometry
journal: Neurocomputing
[…]
[inst1]organization=Luddy School of Informatics, Computing, and Engineering, Indiana University, city=Bloomington, state=IN, country=USA
[…]
[inst2]organization=Department of Mathematics and Computer Science, Drake University,city=Des Moines, state=IA, country=USA
[…]
[inst3]organization=Electronics and Telecommunications Research Institute (ETRI), city=Daejeon, country=Republic of Korea
Figure 1: Our multi-view, multi-scale geometric consistency process. It explicitly models the geometric consistency of the estimated depth map across multiple source views during training at coarse, intermediate, and refine stages. At each stage, the ref. depth estimate map undergo forward-backward-reprojection (see Alg. 2) across NN src. views to generate the GC-penalty map to penalize per-pixel depth error. This approach enables the model to learn geometric constraints more quickly and accurately, leading to improved reconstruction quality during inference.
@@ 1 Introduction
Traditional multi-view stereo (MVS) methods, such as Gipuma [1], Furu [2], COLMAP [3], and Tola [4], primarily focused on addressing photometric and geometric consistency constraints across multiple views. In contrast, recent machine learning-based MVS approaches [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] have used deep networks to extract feature maps and combine them into 3D cost volumes, capturing subtle similarities between features. These advances—ranging from multi-level feature extraction [5, 10] and attention-driven feature matching [5, 19] to refined cost-volume creation [5, 6, 8, 19] and innovative loss formulations [5, 6, 13]—have significantly improved the fidelity of depth estimates and point cloud reconstructions.
@@
This formulation of the loss function offers richer geometric cues that significantly expedite model learning. Our extensive experiments reveal that GC-MVSNet++ nearly halves the number of training iterations compared to other contemporary models [5, 6, 10, 13, 15]. Our approach not only sets a new benchmark for accuracy on the BlendedMVS [20] dataset, but also secures the second position on the DTU [21] and Tanks and Temples [22] advanced benchmark. GC-MVSNet++ is novel in its use of multi-view, multi-scale geometric consistency checks during training with a novel cost regularization network. Extensive ablation experiments further underscore the efficacy of our proposed method. To summarize:
@@
    The versatility of GC-module allows it to be seamlessly integrated into various MVS pipelines to boost geometric consistency during training.
@@ 2 Related Work
Furukawa and Ponce [2] propose a taxonomy that classifies MVS methods into four primary scene representations: volumetric fields [23, 24, 25, 26], point clouds [9, 27], 3D meshes [28], and depth maps [1, 3, 5, 6, 10, 12, 14, 15, 16, 29, 18, 30]. Depth map-based methods can be further divided into traditional techniques that rely on feature detection and geometric constraint solving [1, 2, 3, 29], and learning-based approaches [5, 6, 10, 12, 14, 15, 16], which have gained significant popularity in recent years.
@@
While these methods enhance multi-stage MVS pipelines by refining specific components, they generally lack explicit integration of multi-view geometric cues during the learning process. Consequently, these models depend on the limited geometric information from multiple source views and the cost function formulation during training. Xu and Tao [14] address this limitation with a multi-scale geometric consistency-guided MVS approach that uses multi-hypothesis joint view selection to leverage structured region information for improved candidate hypothesis sampling. They argue that upsampled depth maps from source images can impose geometric constraints on estimates and utilize reprojection error [3, 36, 37] to assess consistency. Unsupervised methods [38, 39] use cross-view consistency to learn geometric features. In contrast, our method employs forward-backward reprojection across multiple source views to directly evaluate and enforce geometric consistency in depth estimates, generating per-pixel penalties for geometrically inconsistent pixels.
@@
  Output: per_pixel_penaltyper\_pixel\_penalty
  Require $M\geq N$
@@
  $D\leftarrow{D_{1}^{gt},...D_{M}^{gt}}$
  $c\leftarrow{c_{1}^{gt},...c_{M}^{gt}}$
@@
   $D^{\prime\prime}_{P^{\prime\prime}_{0}},P^{\prime\prime}_{0}\leftarrow FBR(D_{0},c_{0},D_{i}^{gt},c_{i}^{gt})$ {Alg. 2}
[…]
@@
[…]
Algorithm 1 Geometric Consistency Check Algorithm
@@
  $D_{(R\rightarrow S)}\leftarrow K_{S}\cdot E_{S}\cdot E_{R}^{-1}\cdot K_{R}^{-1}\cdot D_{0}$ {Project}
@@
  $D_{S_{remap}}\leftarrow REMAP(D_{i}^{gt},X_{D_{(R\rightarrow S)}},Y_{D_{(R\rightarrow S)}})$ {Remap}
  $D^{\prime\prime}_{P^{\prime\prime}_{0}}\leftarrow K_{R}\cdot E_{R}\cdot E_{S}^{-1}\cdot K_{S}^{-1}\cdot D_{S_{remap}}$ {Back project}
@@
Algorithm 2 Forward Backward Reprojection (FBR)
3 Methodology
MVS methods take $N$ views as input, including a reference image $I_{0}\in\mathbb{R}^{H\times W\times 3}$ and its paired (N1N-1) source view images $I^{N-1}_{i=1}$, along with the corresponding camera parameters $c_{0},...,c_{N}$, and then to estimate the reference view depth map ($D_{0}$) as the output.
3.1 Network Overview
The architecture of our approach Geometric Consistency MVSNet++, abbreviated as GC-MVSNet++, is shown in Fig. 2. We use a deformable convolution-based [47] feature pyramid network (FPN) [48] architecture (Sec. 3.4) to extract features from input images in a coarse-to-fine manner in three stages. At only the coarse stage, we apply a Global Feature Aggregation (GFA) module with linear attention [35, 5] to leverage global context information within and between reference and source view features. At each stage, we build a correlation-based cost volume of shape $H^{\prime}\times W^{\prime}\times D^{\prime}_{h}\times 1$ using feature maps of shape $N\times H^{\prime}\times W^{\prime}\times C$, where $H^{\prime}$, $W^{\prime}$, and $C$ denote the height, width, and number of channels of a given stage, and $D^{\prime}_{h}$ is the number of depth hypotheses at the corresponding stage. The cost volume is regularized with the proposed dense-connected CostRegNet. We use a winner-takes-all strategy to estimate the depth map $D_{0}$ at each stage.
We employ the GC module at each stage. It checks the geometric consistency of each pixel in $D_{0}$ across $M$ source views and generates $\xi_{p}$ (Sec. 3.2), a pixel-wise factor that is multiplied with the per-pixel depth error ($\xi_{d}$), calculated using a cross-entropy function. It penalizes each pixel in $D_{0}$ for its inconsistency across $M$ source views to accelerate geometric cues learning during training. TransMVSNet [5] (without feature matching transformer and adaptive receptive field module) trained with cross-entropy loss is our baseline (GC-MVS-base), see Table 7.
@@ 3.1 Network Overview
3.2 Multi-View Geometric Consistency Module
GC-MVSNet++ estimates reference depth maps at three stages with different resolutions. At each stage, the GC module takes $D_{0}$, $M$ source view ground truths $D^{gt}_{1},...D^{gt}_{M}$, and their camera parameters $c_{0},...c_{M}$ as input (see Alg. 1). The GC module is then initialized with a geometric inconsistency mask sum (or mask_summask\_sum) of zero at each stage. This mask sum accumulates the inconsistency of each pixel across the $M$ source views. For each source view, the GC module performs forward-backward reprojection of $D_{0}$ to generate the penalty and then adds it to the mask sum.
For each stage, we generate two binary masks of inconsistent pixels, $PDE_{mask}$ and $RDD_{mask}$, by applying thresholds $D_{pixel}$ and $D_{depth}$, and then take a logical-OR of the two to produce a single mask of inconsistent pixels. These inconsistent pixels are assigned a value $1$ and all other pixels, including the consistent and the out-of-scope pixels, are assigned 0 to form a penalty mask. This penalty mask is then added to the mask sum (Alg. 1), which accumulates the penalty mask for each of the $M$ source views to generate a final mask sum with values $\in[0,M]$. Each pixel value indicates the number of inconsistencies of the pixel across the $M$ source views.
From this mask sum, we then generate the inconsistency penalty $\xi_{p}$ for each pixel. Our initial approach generated $\xi_{p}$ by dividing the mask sum by $M$ to normalize within the $[0,1]$. However, we found that using $\xi_{p}$ itself for element-wise multiplication reduces the contribution of perfectly consistent (zero inconsistency) pixels to zero, preventing further improvement of such pixels. To avoid this, we add $1$ so that elements of $\xi_{p}\in[1,2]$. A reference view binary mask is applied on initial $\xi_{p}$ to generate the final $\xi_{p}$, as shown in Fig. 3.
Occlusion and its impact: In multi-view stereo, occluded pixels naturally occur as 3D points often remain invisible from certain views. These hidden pixels significantly affect geometric constraints, leading to the penalization of reference view pixels corresponding to occluded 3D points as inconsistent. To prevent these occluded pixels from disproportionately impacting geometric consistency losses, careful management is essential.
@@ 3.3.1 Dense Connectivity
Let us assume $H_{k}(.)$ denotes the feed-forward 2D convolution operation in a traditional 2D convolutional neural network (CNN), which feeds the output feature map $x_{k-1}$ of the $(k-1)^{th}$ layer to the input of the next layer, $x_{k}=H_{k-1}(x_{k-1})$. To bypass the non-linear transformation, ResNet [53] additionally adds skip-connections, $x_{k}=H_{k-1}(x_{k-1})+x_{k-1}$. DenseNet [52] further improves the information flow between layers by adding a denser connectivity pattern with direct connections from any layer to all its subsequent layers. More precisely, the $k^{th}$ layer receives the concatenation of feature maps $x_{0},x_{1},...x_{k-1}$] from all its preceding layers,
@@ 3.3.3 Decoder Dense Blocks
$x_{out}=concat\left[x^{\prime}_{0},x^{\prime}_{1},...,x^{\prime}_{i-1}\right],$ (2)
@@ 3.4 Deformable Feature Extraction Network
Unlike most MVS methods [5, 6, 10, 13, 57, 58] that use batch normalization [59] during training, we use group normalization throughout our network. As observed in [59], batch normalization provides more consistent and stable training with large batch sizes, but it is inconsistent and has a degrading effect on training with smaller batches. MVS methods are restricted to very small batch sizes, often between 141-4, due to large memory requirements by the 3D-regularization network. Thus, we replaced batch normalization with group normalization layers [54] of group size $4$ across the network. Group normalization performs normalization across a number of channels that is independent of the number of examples in a batch [54]. We also implement weight standardization [55] for all layers (including 3D-convolutional layers) in the network. With these modifications, we achieve stable and reproducible training.
@@ 3.6 Cost Function
3.7 Depth Filtering
To mitigate the impact of inconsistent depth values, we apply the same GC module (refer to Alg. 1) to filter accurate depth maps, with a minor adjustment. Specifically, we substitute the estimated reference depth map ($D_{0}$) with its ground truth counterpart ($D_{0}^{gt}$) in Alg. 1. The hyperparameters are set to $D_{pixel}$=2, $D_{depth}$=0.25, and MM=8, ensuring that only the most erroneous depth values are filtered out. Fig. 6 displays the original and filtered depth maps from various viewpoints of a sample scene in the DTU dataset. The left half of the figure illustrates viewpoints with the most filtered-out values, while the right half shows viewpoints with the fewest. Although we initially identified this depth map rendering issue in the DTU dataset, the approach is broadly applicable to other datasets employing similar methods for generating ground truth depth maps. For the BlendedMVS dataset, we use $D_{pixel}$=0.5, $D_{depth}$=0.05, and MM=10 in the depth filtering process.
Table 1: Quantitative results on DTU and BlendedMVS. Accuracy (Acc), completeness (comp) and overall are in mmmm. We follow Darmon et al. [64] for BlendedMVS evaluation. Bold and underline represents first and second place, respectively. Error maps are shown in Appendix H.